Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

fire four shots

Expert replies
by sanju09 » Fri Mar 19, 2010 5:31 am
An anti aircraft gun can fire four shots at a time. If the probabilities of the first, second, third and the last shot hitting the enemy aircraft are 0.7, 0.6, 0.5 and 0.4, what is the probability that four shots aimed at an enemy aircraft will bring the aircraft down?
(A) 0.036
(B) 0.084
(C) 0.36
(D) 0.64
(E) 0.964
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion
Source: — Problem Solving |

by harshavardhanc » Fri Mar 19, 2010 5:37 am
sanju09 wrote:An anti aircraft gun can fire four shots at a time. If the probabilities of the first, second, third and the last shot hitting the enemy aircraft are 0.7, 0.6, 0.5 and 0.4, what is the probability that four shots aimed at an enemy aircraft will bring the aircraft down?
(A) 0.036
(B) 0.084
(C) 0.36
(D) 0.64
(E) 0.964
I used the complement method here.

prob of guns not hitting the target = .3, .4, .5 ,and .6

Prob that none of the guns hit = .3 * .4 * .5 * .6 = .036

probability that at least one hits = 1- .036 = .084

Am I correct? :(
Regards,
Harsha
Join the discussion

by sanju09 » Fri Mar 19, 2010 5:52 am
harshavardhanc wrote:
sanju09 wrote:An anti aircraft gun can fire four shots at a time. If the probabilities of the first, second, third and the last shot hitting the enemy aircraft are 0.7, 0.6, 0.5 and 0.4, what is the probability that four shots aimed at an enemy aircraft will bring the aircraft down?
(A) 0.036
(B) 0.084
(C) 0.36
(D) 0.64
(E) 0.964
I used the complement method here.

prob of guns not hitting the target = .3, .4, .5 ,and .6

Prob that none of the guns hit = .3 * .4 * .5 * .6 = .036

probability that at least one hits = 1- .036 = .084

Am I correct?
:(
not at subtraction at least, I must say
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by kstv » Fri Mar 19, 2010 6:03 am
Prob of getting a hit
1st shot = 0.7
2nd shot = 0.3 * 0.6=0.8 ( Prob of 1st a miss is 1-0.7)
3rd shot = 0.3*0.4*0.5 = 0.06
4th shot = 0.3*0.4*0.5*0.4= .024
Tot 0.964
Join the discussion

by sanju09 » Fri Mar 19, 2010 6:11 am
kstv wrote:Prob of getting a hit
1st shot = 0.7
2nd shot = 0.3 * 0.6=0.8 ( Prob of 1st a miss is 1-0.7)
3rd shot = 0.3*0.4*0.5 = 0.06
4th shot = 0.3*0.4*0.5*0.4= .024
Tot 0.964
What are we supposed to conclude now?

Is it 1 - 0.024 = 0.964, according to you, or what is your real answer?
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by harshavardhanc » Fri Mar 19, 2010 6:13 am
sanju09 wrote:An anti aircraft gun can fire four shots at a time. If the probabilities of the first, second, third and the last shot hitting the enemy aircraft are 0.7, 0.6, 0.5 and 0.4, what is the probability that four shots aimed at an enemy aircraft will bring the aircraft down?
sanju,


The first sentence conveys that all the four shots are like a burst, i.e all at a time.
An anti aircraft gun can fire four shots at a time
however, the second sentence says that :
the first, second, third and the last shot hitting the enemy
the shots are fired one after the other. There is a difference, which can give rise to two different answers.

Can you clarify or make the question less ambiguous ?
Regards,
Harsha
Join the discussion

by sanju09 » Fri Mar 19, 2010 6:38 am
harshavardhanc wrote:
sanju09 wrote:An anti aircraft gun can fire four shots at a time. If the probabilities of the first, second, third and the last shot hitting the enemy aircraft are 0.7, 0.6, 0.5 and 0.4, what is the probability that four shots aimed at an enemy aircraft will bring the aircraft down?
sanju,


The first sentence conveys that all the four shots are like a burst, i.e all at a time.
An anti aircraft gun can fire four shots at a time
however, the second sentence says that :
the first, second, third and the last shot hitting the enemy
the shots are fired one after the other. There is a difference, which can give rise to two different answers.

Can you clarify or make the question less ambiguous ?
You raised a very good point here. Imagine an anti aircraft from a very good distance if you could. Take case 1 as if the four shots were fired from the single muzzled gun, it would be bound to come out one after the other from the muzzle, so we can number it, fine. Now, take case 2 as if the four shots came out simultaneously from the anti aircraft, in this case it cannot come out from the single muzzle, it would need four different muzzles, directed slightly differently, oh DIFFERENT!! We can still number those, so the answer remains.

Do you know MUZZLE?

Now correct your subtraction work to get to the OA for this question, rest is fine.
The mind is everything. What you think you become. –Lord Buddha



Sanjeev K Saxena
Quantitative Instructor
The Princeton Review - Manya Abroad
Lucknow-226001

www.manyagroup.com
Join the discussion

by harshavardhanc » Fri Mar 19, 2010 7:39 am
sanju09 wrote: correct your subtraction work to get to the OA for this question, rest is fine.
Oooops!!! what a careless mistake!!!!! shame on me!!! :(


.964 :)
Last edited by harshavardhanc on Fri Mar 19, 2010 8:33 am, edited 2 times in total.
Regards,
Harsha
Join the discussion

by RJ43 » Fri Mar 19, 2010 8:18 am
Assuming we only need 1 shot to land for it to go down:

Probability all shots miss = .3 * .4 * .5 *.6 = .0360, or 3.6%

Therefore, the probability of at least 1 shot landing is 1-.036 = 0.964 E.

Yes?
Join the discussion

by hooliganpete » Fri Mar 19, 2010 2:04 pm
Assuming we only need 1 shot to land for it to go down:
Given how detailed we have to be to answer many questions, I would hope a question like this one would specify how many shots are needed to "bring an aircraft down".
Join the discussion